A beaker with 1.00×102 mL of an acetic acid buffer with a pH of 5.000 is sitting on a benchtop. The total molarity of acid and conjugate base in this buffer is 0.100 mol L−1. A student adds 5.80 mL of a 0.440 mol L−1 HCl solution to the beaker. How much will the pH change? The pKa of acetic acid is 4.760.
Answer #1

Given that pH = 5.000

Henderson-Hasselbalch equation:
pH = pKa + log[CH3COO-] / [HC2H3O2- ]
5.00 = 4.760 + log[C2H3O2-] / [HC2H3O2]
0.240 = log[CH3COO-] / [HC2H3O2]
[CH3COO-]/ [HC2H3O2] =10^0.240

[CH3COO-]/ [HC2H3O2] =1.74

1.00×102 mL mL buffer (0.100 M) = 10.0 mmols of Acid + Base        (A + B from now on)

[CH3COO-]/ [HC2H3O2] =1.74

[CH3COO-] = [HC2H3O2] 1.74 ----1

10.0 mmol = [CH3COO-] + [HC2H3O2] ---2

10.0 mmol =[HC2H3O2] 1.74+ [HC2H3O2]

10.0 mmol =[HC2H3O2] 2.74

[HC2H3O2] = 3.65 mmol = acetic acid

Therefore

[CH3COO-] = [HC2H3O2] 1.74 ----1

[CH3COO-] = 3.65* 1.74 = 6.35 mmol

Now that we have our mmol of Acid and acetate, we can see what remains after HCl is added:

5.80 mL HCl (0.440 M) = 2.552 mmol HCl added

            Acetate         +           HCl        --->    Acetic acid            +             H2O

Before            6.35                      2.552                3.65

Change          -2.552                      -2.552                +2.552

Final              3.798                        0                      6.202

We still have a buffer, so we can use pH = pKa + log[CH3COO-] / [HC2H3O2- ] again to find the pH:

pH = 4.760 + log (3.798/6.202)

pH = 4.760 + (0..213)

= 4.547

ΔpH: 4.547 - 5.000 = -0.453